# UniMoT: Unified Molecule-Text Language Model with Discrete Token Representation

### v1

@@ Appendix A Details of Causal Q-Former
The Q-Former operates as a query-based transformer that utilizes learnable query vectors to interact with molecule features extracted by a frozen encoder. These queries are essential for extracting relevant information from the molecule features. The Q-Former comprises both a molecule transformer and a text transformer, sharing self-attention layers. The molecule transformer incorporates cross-attention layers between self-attention and feed-forward layers, while the text transformer architecture is based on BERT [9]. Q-Former employs a cross-attention mechanism where the query vectors selectively attend to different aspects of the molecule features, allowing the model to capture critical details necessary for understanding and generating textual descriptions of molecular properties.
@@
$\mathcal{L}_{\text{MTC}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp((\boldsymbol{z}_{M}^{i})^{T}\boldsymbol{y}^{i}/\tau)}{\sum_{j=1}^{B}\exp((\boldsymbol{z}_{M}^{i})^{T}\boldsymbol{y}^{j}/\tau)}-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp((\boldsymbol{y}^{i})^{T}\boldsymbol{z}_{M}^{i}/\tau)}{\sum_{j=1}^{B}\exp((\boldsymbol{y}^{i})^{T}\boldsymbol{z}_{M}^{j}/\tau)},$ (4)
@@
$\mathcal{L}_{\text{MTM}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\phi(\boldsymbol{z}_{M}\mid\mathbf{X}^{i},\boldsymbol{t}^{i}))}{\sum_{j=1}^{B}\exp(\phi(\boldsymbol{z}_{M}\mid\mathbf{X}^{i},\boldsymbol{t}^{j}))+\sum_{j=1}^{B}\exp(\phi(\boldsymbol{z}_{M}\mid\mathbf{X}^{j},\boldsymbol{t}^{i}))},$ (5)
@@ Appendix B Details of Datasets
PubChem Captioning, Retrieval, Generation Molecule Captioning, Molecule-Text Retrieval, Caption-guided Molecule Generation Generates descriptions and retrieves text/molecules based on input molecules/text, and guides molecule generation from captions. PubChem URL Public Domain
@@
  • BBBP [53]: The Blood-Brain Barrier Penetration dataset predicts the ability of molecules to penetrate the blood-brain barrier.
  • Tox21 [53]: This dataset is part of the Toxicology in the 21st Century initiative, used for toxicity prediction.
  • ToxCast [53]: Another toxicity prediction dataset with a broader range of biological assays.
  • Sider [53]: Side Effect Resource database, used for predicting drug side effects.
  • ClinTox [53]: Clinical Toxicity dataset for predicting clinical trial toxicity outcomes.
  • MUV [53]: Maximum Unbiased Validation dataset for virtual screening.
  • HIV [53]: Human Immunodeficiency Virus dataset for predicting anti-HIV activities.
  • BACE [53]: Beta-Secretase 1 dataset for predicting inhibitors of the BACE-1 enzyme, relevant for Alzheimer’s research.
  • QM9 [12]: The quantum mechanics properties dataset, where the objective is to predict key quantum mechanics properties of a given molecule, such as HUMO, LUMO, and the HUMO-LUMO gap.
@@
  • PubChem [18]: A large dataset of chemical molecules used for generating textual descriptions of molecular structures.
  • ChEBI-20 [11]: A subset of the Chemical Entities of Biological Interest database, provides structured and detailed descriptions of molecules.
@@
  • PubChem [18]: Used for both molecule-to-text (M2T) and text-to-molecule (T2M) retrieval tasks.
  • PCdes [58]: Another dataset for evaluating M2T and T2M retrieval accuracy.
  • MoMu [39]: Dataset specifically designed for molecule-text interactions and retrieval tasks.
@@
  • Mol-Instructions [12]: These datasets include tasks such as caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. They are used to evaluate the model’s ability to generate molecular structures based on textual descriptions and other related tasks.
  • PubChem [18]: Used for caption-guided molecule generation, generating molecular structures based on textual descriptions.
  • USPTO [12]: Used for reagent prediction, forward reaction prediction, and retrosynthesis, providing data for predicting reagents, reaction outcomes, and retrosynthetic pathways.
@@ Appendix C Details of Training
During Stage-1, we only connect the molecule encoder and the Causal Q-Former, leaving out other blocks. We leverage the pretrained molecule encoder from MoleculeSTM [31], which has undergone extensive contrastive learning with molecule-text pairs. We utilize the PubChem [18] dataset for pretraining, keeping the molecule encoder frozen while updating only the Causal Q-Former. Both queries and text serve as input to the Causal Q-Former, while only queries serve as input in subsequent stages. Inspired by BLIP-2 [23], we employ three tailored objectives – Molecule-Text Contrastive Learning (MTC), Molecule-Text Matching (MTM), and Molecule-grounded Text Generation (MTG) – for the pretraining of the Causal Q-Former, as detailed in Appendix A.
@@
We connect the Causal Q-Former with the subsequent blocks and train the molecule tokenizer using the objective defined in Equation (2). Following the approach of RetMol [49], we utilize SMILES strings [50] to represent molecules, and employ the pretrained ChemFormer [17] as the generative model. Specifically, we leverage the SMILES encoder and SMILES decoder components provided by ChemFormer. We utilize PubChem [18] and CheBI-20 [11] datasets, keeping the molecule encoder, SMILES encoder, and SMILES decoder frozen, while updating the Causal Q-Former, codebook, and adapter. Once optimized, the molecule tokenizer remains unchanged throughout the subsequent stages.
@@
We connect the molecule tokenizer with the LLM and employ the LM objective defined in Equation (3) to pretrain the LLM. We utilize Llama [44] as the default LLM. To construct the unified molecule-text vocabulary, we merge 2048 molecule codes with the original text vocabulary. Pretraining the LLM involves molecule-to-text autoregression and text-to-molecule autoregression, aimed at enhancing UniMoT’s multi-modal comprehension and generation capabilities. We utilize datasets PubChem [18] and CheBI-20 [11] for this purpose. To enhance efficiency, we train the LLM using LoRA tuning [14].
@@
We perform instruction tuning to align UniMoT with human instructions through supervised fine-tuning on seven tasks: molecular property prediction, molecule captioning, molecule-text retrieval, caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. For the molecular property prediction task, we utilize the quantum mechanics properties dataset [12] for regression prediction and the MoleculeNet [53] datasets for property classification. For the molecule captioning and molecule-text retrieval tasks, we employ datasets PubChem [18], PCdes [58], and MoMu [39]. For the molecule generation tasks, we utilize the Mol-Instructions [12] datasets to conduct instruction tuning. We fine-tune UniMoT for 10 epochs on each task using the same optimizer, learning rate scheduler, and LoRA configurations as in Stage-3 pretraining. Instruction samples for comprehension and generation tasks are shown in Table 8.
@@ Appendix D Details and More Results of Experiments
Property prediction aims to anticipate a molecule’s intrinsic physical and chemical properties based on its structural or sequential characteristics. In the regression task, we conduct experiments on the quantum mechanics properties dataset QM9 [12], where the objective is to predict key quantum mechanics properties of a given molecule, such as HUMO, LUMO, and the HUMO-LUMO gap. We compare UniMoT against several baselines, including Alpaca [42], Baize [55], Llama-2-7B [44], Vicuna-13B [7], Mol-Instructions [12], and InstructMol [6]. Mean Absolute Error (MAE) serves as our evaluation metric. The performance of the regression task on the QM9 dataset is presented in Table 10. Compared to previous single-modal instruction-tuned LLMs and molecular LLMs, UniMoT exhibits further improvement on the regression task, showcasing its fundamental comprehension abilities in molecular contexts.
Table 10: Mean Absolute Error (MAE) of molecular property prediction task (regression) on the QM9 [12] dataset. Bold indicates the best performance and underline indicates the second best performance. $\Delta\epsilon$ is the HOMO-LUMO energy gap.
@@
Alpaca (Llama-7B) [42] - - - 322.109
Baize (Llama-7B) [55] - - - 261.343
Llama-2-7B [44] 0.7367 0.8641 0.5152 0.7510
Vicuna-13B [7] 0.7135 3.6807 1.5407 1.9783
Mol-Instructions (Llama-7B) [12] 0.0210 0.0210 0.0203 0.0210
InstructMol (Vicuna-7B) [6] 0.0048 0.0050 0.0061 0.0050
@@
The molecule captioning task involves generating a comprehensive description of a molecule. For this task, we compare UniMoT with several baselines: MolT5 [11], MoMu [39], InstructMol [6], MolCA [33], and 3D-MoLM [25]. We adopt BLEU [37], ROUGE [27], and METEOR [1] as the evaluation metrics. The performance of UniMoT in the molecule captioning task on the CheBI-20 [11] dataset is presented in Table 11. Some concrete examples of molecule captioning task are presented in Table 12. From the results, it is evident that UniMoT consistently outperforms the baselines by a significant margin. These results underscore the effectiveness of the molecule tokenizer in providing molecule tokens with high-level molecular and textual information, thus enhancing molecule comprehension.
Table 11: Performance (%) of molecule captioning task on the CheBI-20 [11] dataset. Bold indicates the best performance and underline indicates the second best performance.
@@
T5-Small [38] 50.1 41.5 60.2 44.6 54.5 53.2
T5-Base [38] 51.1 42.3 60.7 45.1 55.0 53.9
T5-Large [38] 55.8 46.7 63.0 47.8 56.9 58.6
MolT5-Small (T5-Small) [11] 51.9 43.6 62.0 46.9 56.3 55.1
MolT5-Base (T5-Base) [11] 54.0 45.7 63.4 48.5 57.8 56.9
MolT5-Large (T5-Large) [11] 59.4 50.8 65.4 51.0 59.4 61.4
MoMu-Small (T5-Small) [39] 53.2 44.5 - - 56.4 55.7
MoMu-Base (T5-Base) [39] 54.9 46.2 - - 57.5 57.6
MoMu-Large (T5-Large) [39] 59.9 51.5 - - 59.3 59.7
InstructMol (Vicuna-7B) [6] 47.5 37.1 56.6 39.4 50.2 50.9
MolCA (OPT-125M) [33] 61.6 52.9 67.4 53.3 61.5 63.9
MolCA (OPT-1.3B) [33] 63.9 55.5 69.7 55.8 63.6 66.9
@@
The molecule-text retrieval task involves using a molecule to retrieve text (M2T) and using text to retrieve a molecule (T2M). We compare UniMoT with several baselines: Sci-BERT [2], KV-PLM [58], MoMu [39], MoleculeSTM [31], MolCA [33], and 3D-MoLM [25]. We report the performance of retrieval using a batch of 64 random samples and the entire test set, evaluated with the metrics of Accuracy and Recall@20. We use the checkpoint from Stage-1 of pretraining. Performance on the PCdes [58] and MoMu [39] datasets is shown in Table 13. UniMoT demonstrates superior performance over the baselines on molecule-text retrieval, particularly in molecule-to-text retrieval. This demonstrates that UniMoT has learned fine-grained alignment between molecules and text, and it can understand molecule-text interactions through the introduction of the Causal Q-Former.
Table 13: Accuracy (%) of molecule-text retrieval task on the PCdes [58] and MoMu [39] datasets. Bold indicates the best performance and underline indicates the second best performance. We report the performance of retrieval using a batch of 64 random samples and the entire test set.
(a) Accuracy (%) of molecule-text retrieval task on the PCdes [58] dataset.
@@
Sci-BERT [2] 62.6 61.8 60.7 60.8
KV-PLM [58] 77.9 65.0 75.9 64.3
MoMu (Sci-BERT) [39] 80.6 77.0 79.1 75.5
MoMu (KV-PLM) [39] 81.1 80.2 80.2 79.0
MoleculeSTM [31] 86.2 83.9 84.6 85.1
MolCA (OPT-1.3B) [33] 91.4 88.4 90.5 87.6
3D-MoLM (Llama-2-7B) [25] 92.3 89.6 91.2 88.5
UniMoT (Llama-2-7B) 92.6 89.4 91.6 88.3
(b) Accuracy (%) of molecule-text retrieval task on the MoMu [39] dataset.
@@
Sci-BERT [2] 1.4 1.6 0.3 0.3
KV-PLM [58] 1.5 1.3 0.5 0.3
MoMu (Sci-BERT) [39] 45.7 40.0 43.3 43.4
MoMu (KV-PLM) [39] 46.2 38.5 43.7 43.5
MoleculeSTM [31] 81.8 81.9 75.8 74.5
MolCA (OPT-1.3B) [33] 83.7 84.3 88.6 87.3
3D-MoLM (Llama-2-7B) [25] 84.9 85.4 89.9 88.7

### v2

@@ Appendix A Details of Causal Q-Former
The Q-Former operates as a query-based transformer that utilizes learnable query vectors to interact with molecule features extracted by a frozen encoder. These queries are essential for extracting relevant information from the molecule features. The Q-Former comprises both a molecule transformer and a text transformer, sharing self-attention layers. The molecule transformer incorporates cross-attention layers between self-attention and feed-forward layers, while the text transformer architecture is based on BERT Devlin et al. [2018]. Q-Former employs a cross-attention mechanism where the query vectors selectively attend to different aspects of the molecule features, allowing the model to capture critical details necessary for understanding and generating textual descriptions of molecular properties.
@@
$\mathcal{L}_{\text{MTC}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp((\boldsymbol{z}_{M}^{i})^{T}\boldsymbol{y}^{i}/\tau)}{\sum_{j=1}^{B}\exp((\boldsymbol{z}_{M}^{i})^{T}\boldsymbol{y}^{j}/\tau)}\\ -\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp((\boldsymbol{y}^{i})^{T}\boldsymbol{z}_{M}^{i}/\tau)}{\sum_{j=1}^{B}\exp((\boldsymbol{y}^{i})^{T}\boldsymbol{z}_{M}^{j}/\tau)},$ (4)
@@
$\mathcal{L}_{\text{MTM}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\phi(\boldsymbol{z}_{M}\mid\mathbf{X}^{i},\boldsymbol{t}^{i}))}{\sum_{j=1}^{B}\exp(\phi(\boldsymbol{z}_{M}\mid\mathbf{X}^{i},\boldsymbol{t}^{j}))+\sum_{j=1}^{B}\exp(\phi(\boldsymbol{z}_{M}\mid\mathbf{X}^{j},\boldsymbol{t}^{i}))}.$
@@ Appendix B Details of Datasets
PubChem Captioning, Retrieval, Generation Molecule Captioning, Molecule-Text Retrieval, Caption-guided Molecule Generation Generates descriptions, retrieves text / molecules based on input molecules / text, and guides molecule generation from captions. PubChem URL Public Domain
@@
  • BBBP Wu et al. [2018]: The Blood-Brain Barrier Penetration dataset predicts the ability of molecules to penetrate the blood-brain barrier.
  • Tox21 Wu et al. [2018]: This dataset is part of the Toxicology in the 21st Century initiative, used for toxicity prediction.
  • ToxCast Wu et al. [2018]: Another toxicity prediction dataset with a broader range of biological assays.
  • Sider Wu et al. [2018]: Side Effect Resource database, used for predicting drug side effects.
  • ClinTox Wu et al. [2018]: Clinical Toxicity dataset for predicting clinical trial toxicity outcomes.
  • MUV Wu et al. [2018]: Maximum Unbiased Validation dataset for virtual screening.
  • HIV Wu et al. [2018]: Human Immunodeficiency Virus dataset for predicting anti-HIV activities.
  • BACE Wu et al. [2018]: Beta-Secretase 1 dataset for predicting inhibitors of the BACE-1 enzyme, relevant for Alzheimer’s research.
  • QM9 Fang et al. [2023]: The quantum mechanics properties dataset, where the objective is to predict key quantum mechanics properties of a given molecule, such as HUMO, LUMO, and the HUMO-LUMO gap.
@@
  • PubChem Kim et al. [2023]: A large dataset of chemical molecules used for generating textual descriptions of molecular structures.
  • ChEBI-20 Edwards et al. [2022]: A subset of the Chemical Entities of Biological Interest database, provides structured and detailed descriptions of molecules.
@@
  • PubChem Kim et al. [2023]: Used for both molecule-to-text (M2T) and text-to-molecule (T2M) retrieval tasks.
  • PCdes Zeng et al. [2022]: Another dataset for evaluating M2T and T2M retrieval accuracy.
  • MoMu Su et al. [2022]: Dataset specifically designed for molecule-text interactions and retrieval tasks.
@@
  • Mol-Instructions Fang et al. [2023]: This benchmark includes tasks such as caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. It is used to evaluate the model’s ability to generate molecular structures based on textual descriptions and other related tasks.
  • PubChem Kim et al. [2023]: Used for caption-guided molecule generation, generating molecular structures based on textual descriptions.
  • USPTO Fang et al. [2023]: Used for reagent prediction, forward reaction prediction, and retrosynthesis, providing data for predicting reagents, reaction outcomes, and retrosynthetic pathways.
@@ Appendix C Details of Training
During Stage-1, we only connect the molecule encoder and the Causal Q-Former, leaving out other blocks. We leverage the pretrained molecule encoder from MoleculeSTM Liu et al. [2023a], which has undergone extensive contrastive learning with molecule-text pairs. We utilize the PubChem Kim et al. [2023] dataset for pretraining, keeping the molecule encoder frozen while updating only the Causal Q-Former. Both queries and text serve as input to the Causal Q-Former, while only queries serve as input in subsequent stages. Inspired by BLIP-2 Li et al. [2023], we employ three tailored objectives – Molecule-Text Contrastive Learning (MTC), Molecule-Text Matching (MTM), and Molecule-grounded Text Generation (MTG) – for the pretraining of the Causal Q-Former, as detailed in Appendix A.
@@
We connect the Causal Q-Former with the subsequent blocks and train the molecule tokenizer using the objective defined in Equation (2). We utilize SMILES strings Weininger [1988] to represent molecules, and employ the pretrained ChemFormer Irwin et al. [2022] as the generative model. Specifically, we leverage the SMILES encoder and SMILES decoder components provided by ChemFormer. We utilize PubChem Kim et al. [2023] and ChEBI-20 Edwards et al. [2022] datasets, keeping the molecule encoder, SMILES encoder, and SMILES decoder frozen, while updating the Causal Q-Former, codebook, and adapter. Once optimized, the molecule tokenizer remains unchanged throughout the subsequent stages.
@@
We connect the molecule tokenizer with the LLM and employ the LM objective defined in Equation (3) to pretrain the LLM. We utilize Llama Touvron et al. [2023] as the default LLM. To construct the unified molecule-text vocabulary, we merge 2048 molecule codes with the original text vocabulary. Pretraining the LLM involves molecule-to-text autoregression and text-to-molecule autoregression, aimed at enhancing UniMoT’s multi-modal comprehension and generation capabilities. We utilize datasets PubChem Kim et al. [2023] and ChEBI-20 Edwards et al. [2022] for this purpose. To enhance efficiency, we train the LLM using LoRA Hu et al. [2021].
@@
We perform instruction tuning to align UniMoT with human instructions through supervised fine-tuning on seven tasks: molecular property prediction, molecule captioning, molecule-text retrieval, caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. For the molecular property prediction task, we utilize the quantum mechanics properties dataset Fang et al. [2023] for regression prediction and the MoleculeNet Wu et al. [2018] datasets for property classification. For the molecule captioning and molecule-text retrieval tasks, we employ datasets PubChem Kim et al. [2023], PCdes Zeng et al. [2022], and MoMu Su et al. [2022]. For the molecule generation tasks, we utilize the Mol-Instructions Fang et al. [2023] benchmark to conduct instruction tuning. We fine-tune UniMoT for 10 epochs on each task using the same optimizer, learning rate scheduler, and LoRA configurations as in Stage-3 pretraining. Instruction samples for comprehension and generation tasks are shown in Table 6.
@@ Appendix D Details and More Results of Experiments
Property prediction aims to anticipate a molecule’s intrinsic physical and chemical properties based on its structural or sequential characteristics. In the regression task, we conduct experiments on the quantum mechanics properties dataset QM9 Fang et al. [2023], where the objective is to predict key quantum mechanics properties of a given molecule, such as HUMO, LUMO, and the HUMO-LUMO gap. We compare UniMoT against several baselines, including Alpaca Taori et al. [2023], Baize Xu et al. [2023], Llama-2-7B Touvron et al. [2023], Vicuna-13B Chiang et al. [2023], Mol-Instructions Fang et al. [2023], and InstructMol Cao et al. [2023]. Mean Absolute Error (MAE) serves as our evaluation metric. The performance of the regression task on the QM9 dataset is presented in Table 8. Compared to previous single-modal instruction-tuned LLMs and molecular LLMs, UniMoT exhibits further improvement on the regression task, showcasing its fundamental comprehension abilities in molecular contexts.
@@
Alpaca (Llama-7B) - - - 322.109
Baize (Llama-7B) - - - 261.343
Llama-2-7B 0.7367 0.8641 0.5152 0.7510
Vicuna-13B 0.7135 3.6807 1.5407 1.9783
Mol-Instructions (Llama-7B) 0.0210 0.0210 0.0203 0.0210
InstructMol (Vicuna-7B) 0.0048 0.0050 0.0061 0.0050
@@
Table 8: Mean Absolute Error (MAE) of molecular property prediction task (regression) on the QM9 Fang et al. [2023] dataset. Bold indicates the best performance and underline indicates the second best performance. $\Delta\epsilon$ is the HOMO-LUMO energy gap.
@@
The molecule captioning task involves generating a comprehensive description of a molecule. For this task, we compare UniMoT with several baselines: MolT5 Edwards et al. [2022], MoMu Su et al. [2022], InstructMol Cao et al. [2023], MolCA Liu et al. [2023b], and 3D-MoLM Li et al. [2024]. We adopt BLEU Papineni et al. [2002], ROUGE Lin [2004], and METEOR Banerjee and Lavie [2005] as the evaluation metrics. The performance of UniMoT in the molecule captioning task on the ChEBI-20 Edwards et al. [2022] dataset is presented in Table 9. Some concrete examples of molecule captioning task are presented in Table 10. From the results, it is evident that UniMoT consistently outperforms the baselines by a significant margin. These results underscore the effectiveness of the molecule tokenizer in providing molecule tokens with high-level molecular and textual information, thus enhancing molecule comprehension.
@@
T5-Small 50.1 41.5 60.2 44.6 54.5 53.2
T5-Base 51.1 42.3 60.7 45.1 55.0 53.9
T5-Large 55.8 46.7 63.0 47.8 56.9 58.6
MolT5-Small (T5-Small) 51.9 43.6 62.0 46.9 56.3 55.1
MolT5-Base (T5-Base) 54.0 45.7 63.4 48.5 57.8 56.9
MolT5-Large (T5-Large) 59.4 50.8 65.4 51.0 59.4 61.4
MoMu-Small (T5-Small) 53.2 44.5 - - 56.4 55.7
MoMu-Base (T5-Base) 54.9 46.2 - - 57.5 57.6
MoMu-Large (T5-Large) 59.9 51.5 - - 59.3 59.7
InstructMol (Vicuna-7B) 47.5 37.1 56.6 39.4 50.2 50.9
MolCA (OPT-125M) 61.6 52.9 67.4 53.3 61.5 63.9
MolCA (OPT-1.3B) 63.9 55.5 69.7 55.8 63.6 66.9
@@
Table 9: Performance (%) of molecule captioning task on the ChEBI-20 Edwards et al. [2022] dataset. Bold indicates the best performance and underline indicates the second best performance.
@@
[…]
Sci-BERT 85.3 98.7 84.2 98.4 41.7 87.3 40.2 86.8
KV-PLM 86.1 98.6 85.2 98.5 42.8 88.5 41.7 87.8
MoMu (Sci-BERT) 87.6 99.2 86.4 99.4 47.3 90.8 48.1 89.9
MoMu (KV-PLM) 88.2 99.4 87.3 99.4 48.5 91.6 49.5 90.7
MoleculeSTM 90.5 99.6 88.6 99.5 52.7 92.9 53.2 92.5
MolCA (OPT-1.3B) 92.6 99.8 91.3 99.5 67.9 94.4 68.6 93.3
3D-MoLM (Llama-2-7B) 93.5 100.0 92.9 99.6 69.1 95.9 70.1 94.9
UniMoT (Llama-2-7B) 93.6 100.0 92.7 99.4 69.5 96.3 69.8 94.4
Table 11: Performance (%) of molecule-text retrieval task on the PubChem Kim et al. [2023] dataset. Bold indicates the best performance and underline indicates the second best performance.
@@
Sci-BERT 62.6 61.8 60.7 60.8
KV-PLM 77.9 65.0 75.9 64.3
MoMu (Sci-BERT) 80.6 77.0 79.1 75.5
MoMu (KV-PLM) 81.1 80.2 80.2 79.0
MoleculeSTM 86.2 83.9 84.6 85.1
MolCA (OPT-1.3B) 91.4 88.4 90.5 87.6
3D-MoLM (Llama-2-7B) 92.3 89.6 91.2 88.5
UniMoT (Llama-2-7B) 92.6 89.4 91.6 88.3
(a) Accuracy (%) of molecule-text retrieval task on the PCdes Zeng et al. [2022] dataset.
Model Retrieval in batch Retrieval in test set
M2T (%) T2M (%) M2T (%) T2M (%)
Sci-BERT 1.4 1.6 0.3 0.3
KV-PLM 1.5 1.3 0.5 0.3
MoMu (Sci-BERT) 45.7 40.0 43.3 43.4
MoMu (KV-PLM) 46.2 38.5 43.7 43.5
MoleculeSTM 81.8 81.9 75.8 74.5
MolCA (OPT-1.3B) 83.7 84.3 88.6 87.3
3D-MoLM (Llama-2-7B) 84.9 85.4 89.9 88.7
@@
(b) Accuracy (%) of molecule-text retrieval task on the MoMu Su et al. [2022] dataset.
Table 12: Accuracy (%) of molecule-text retrieval task on the PCdes Zeng et al. [2022] and MoMu Su et al. [2022] datasets. Bold indicates the best performance and underline indicates the second best performance. We report the performance of retrieval using a batch of 64 random samples and the entire test set.
The molecule-text retrieval task involves using a molecule to retrieve text (M2T) and using text to retrieve a molecule (T2M). We compare UniMoT with several baselines: Sci-BERT Beltagy et al. [2019], KV-PLM Zeng et al. [2022], MoMu Su et al. [2022], MoleculeSTM Liu et al. [2023a], MolCA Liu et al. [2023b], and 3D-MoLM Li et al. [2024]. We report the performance of retrieval using a batch of 64 random samples and the entire test set, evaluated with the metrics of Accuracy and Recall@20. We use the checkpoint from Stage-1 of pretraining. UniMoT is evaluated on the datasets of PubChem Kim et al. [2023], PCdes Zeng et al. [2022], and MoMu Su et al. [2022]. Performance on the PubChem dataset is shown in Table 11, while performance on the PCdes Zeng et al. [2022] and MoMu Su et al. [2022] datasets is shown in Table 12. UniMoT demonstrates superior performance over the baselines on molecule-text retrieval, particularly in molecule-to-text retrieval. This demonstrates that UniMoT has learned fine-grained alignment between molecules and text, and it can understand molecule-text interactions through the introduction of the Causal Q-Former.